OpenStudy (anonymous):

Find the derivativate using the chain rule of ((1/2)(3x)^-1/2)

OpenStudy (konradzuse):

\[\frac{1}{2} 3x^{\frac{-1}{2}} ???\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

umm, no need for the chain rule here

OpenStudy (anonymous):

\[\frac{3}{ 2 } \frac{d}{dx}3x^{-\frac{1}{2}}\]